# The NeRFect Match: Exploring NeRF Features for Visual Localization

### v1

@@ 6 Conclusion and Limitations
In this work, we have taken initial steps towards leveraging NeRF as the primary representation for the task of camera localization. To achieve this, we have thoroughly examined the performance of NeRF features in the localization task, considering various architectural designs, feature extraction from different encoder layers, and diverse training configurations. Additionally, we have demonstrated that NeRF can remove the need for the original image set for coarse localization. Our results suggest that NeRF features are highly effective for 2D-3D matching. Specifically, our novel method, NeRFMatch, sets a new state-of-the-art for localization performance on the Cambridge Landmarks dataset.
While NeRFMatch marks a significant step towards comprehensive localization using NeRF, it also highlights several limitations that necessitate further research. Specifically, we observe a noticeable performance gap when applying our method to indoor 7-Scenes dataset.
@@
Appendix A NeRF Implementation Details
To properly train NeRF in such a scenario, we use a pre-trained semantic segmentation model [19] and mask out the sky and transient objects: pedestrians, bicycles, and vehicles. These objects occupy only a minor part of the captured images and are excluded from the loss computation during the training process. Analogous methods for masking in sky regions and/or dynamic object areas have been implemented in other works focused on the reconstruction of urban scenes [61, 47, 70]. We show examples of semantic segmentation in Fig. 5 and its effect on synthesized views in
Appendix B NeRFMatch Implementation Details
@@
top1 top5 (M.Match) top5 (M.3D)
Real Query - Real DB 10.5/0.3/87.1 10.0/0.3/87.5 10.4/0.3/86.7
Real Query - Synthetic DB 11.4/0.3/79.8 10.8/0.3/83.9 11.5/0.3/81.1
Table 6: Retrieval pair ablation on Cambridge Landmarks. We compare NeRFMatch performance using different retrieval pairs that are computed against real database images versus NeRF rendered database images, under different topk used for testing. We report the median pose errors and localization recall averaged over 5 scenes. M.Match stands for merging matches and M.3D stands for merging 3D points.
Training pairs.  We use the same training pairs [^2] Image pairs are available from https://cvg-data.inf.ethz.ch/pixloc_CVPR2021/ generated by PixLoc[53] which were computed based on image covisibility within the training split. During training, for each train image we load its top-$20$ covisible pairs. For each training epoch, we then randomly sample $10000$ training pairs from those covisible pairs for each scene. In the case, we train multiple scenes, we merge those pairs across scenes which allows us to balance the training samples across different scenes.
Appendix C Leveraging Multiple Reference Poses
In this section, we ablate the best way to leverage multiple reference poses to further improve our NeRFMatch model. We propose two options to work with top-$k$ retrieval poses with $k>1$. In the M.Match mode, we simply merge the matched correspondences after exhaustively performing $k$ times matching with NeRFMatch. In the M.3D mode, we first merge the 3D points rendered from $k$ retrieval poses and keep a subset of 3D points that are commonly observed by multiple views. We then random select up to $6003$D points from the subset and apply NeRFMatch on those merged 3D points. As shown in Tab. 6, NeRFMatch benefits from multiple references in M.Match mode, especially for pairs obtained from synthesized training images. However, multiple views sometimes lead to worse performance in M.3D, since useful 3D information is dropped at an early stage and cannot be recovered.

### v2

@@ 6 Conclusion and Limitations
In this work, we have taken initial steps towards leveraging NeRF as the primary representation for the task of camera localization. To achieve this, we have thoroughly examined the performance of NeRF features in the localization task, considering various architectural designs, feature extraction from different encoder layers, and diverse training configurations. Additionally, we have demonstrated that NeRF can remove the need for the original image set for coarse localization. Our results suggest that NeRF features are highly effective for 2D-3D matching.
While NeRFMatch marks a significant step towards comprehensive localization using NeRF, it also highlights several limitations that necessitate further research. Specifically, we observe a noticeable performance gap when applying our method to indoor 7-Scenes dataset. NeRFMatch | Supplementary
In this supplementary document, we provide further details regarding our proposed method and qualitative results. We describe our implementation details for NeRF in Appendix 0.A, and for NeRFMatch in Appendix 0.B. Then, we present additional analysis and discussion of our method in Appendix 0.C.
Appendix 0.A NeRF Implementation Details
To properly train NeRF in such a scenario, we use a pre-trained semantic segmentation model [19] and mask out the sky and transient objects: pedestrians, bicycles, and vehicles. These objects occupy only a minor part of the captured images and are excluded from the loss computation during the training process. Analogous methods for masking in sky regions and/or dynamic object areas have been implemented in other works focused on the reconstruction of urban scenes [61, 47, 70]. We show examples of semantic segmentation in Fig. 5 and its effect on synthesized views in Fig. 6.
@@ Appendix 0.A NeRF Implementation Details
Appendix 0.B NeRFMatch Implementation Details
We summarize average runtime performance for NeRF and both matching models in Tab. 6.
Table 6: Runtime. We show runtime of NeRFMatch-Mini and NeRFMatch. For pose refinement we are using optimization refinement for NeRFMatch-Mini and iterative refinement for NeRFMatch.
NeRF type NeRFMatch-Mini NeRFMatch
NeRF render 141ms 141ms
Image-to-NeRF matching 37ms 157ms
Pose refinement 398ms 141ms
Training pairs.  We use the same training pairs***Image pairs are available from https://cvg-data.inf.ethz.ch/pixloc_CVPR2021/ generated by PixLoc[53] which were computed based on image covisibility within the training split. During training, for each train image we load its top-$20$ covisible pairs. For each training epoch, we then randomly sample $10000$ training pairs from those covisible pairs for each scene. In the case, we train multiple scenes, we merge those pairs across scenes which allows us to balance the training samples across different scenes.
@@ Appendix 0.B NeRFMatch Implementation Details
Optimization refinement.  Similar to iNeRF [73], we are doing a forward pass through frozen NeRF MLP layers using an estimated pose as the initial camera pose. Instead of rendering the entire image, we sample and render $3600$ rays, which are equally spread in a grid structure across the image plane. The we apply a regular photometric loss between the query image and the rendered image and backpropagate to update the initial camera pose. Instead of using the raw updated camera pose, we render the NeRF features and match them with the NeRFMatch to obtain the final camera pose.
Appendix 0.C Additional Details
NeRF backbones.  In this section, we evaluate additional NeRF type - Instant NGP [45] in comparison to MipNeRF [3]. We use MipNeRF for our experiments in the main paper . As shown in Tab. 7, Instant NGP performs significantly worse. We hypothesize that this is due to noisy depth reconstruction that is typical for Instant NGP.
N
Table 7: NeRF backbone ablation on Cambridge Landmarks. We compare NeRFMatch-Mini and NeRFMatch performances using Instant NGP.
[…]
NeRFMatch-Mini NeRFMatch
Instant NGP 41.1/0.7/44.4 28.1/0.5/61.3
MipNeRF 20.0/0.4/69.7 13.3/0.3/80.8
Impact of scene sizes.  Scene size affects both NeRF and localization performance, often coupled with scene content and camera pose distribution. Ranking scenes by localization errors (lower is better) leads to OldHospital ($50\times 40m^{2}$) $>$ KingsCollege ($140\times 40m^{2}$) $>$ ShopFacade ($35\times 25m^{2}$) for outdoor and stairs ($2.5\times 2\times 1.5m^{3}$) $>$ pumpkin ($2.5\times 2\times 1m^{3}$) $>$ redkitchen ($4\times 4\times 1.5m^{3}$) $>$ chess ($3\times 2\times 1m^{3}$) for indoor. This suggests that smaller scenes (OldHospital, stairs) can be more challenging than larger scenes (KingsCollege, redkitchen) due to challenging contents like repetitive structures and texture-less regions.
Image retrieval on synthesized views.  The goal of NeRF-only experiment is to verify the possibility to use NeRF as the only scene representation removing the need to maintain the original image collection. Our experiments show a slight performance decrease due to the domain gap between rendered and real images. Yet, we did not claim an efficient solution for online image retrieval and NeRF rendering. Future research is needed to improve its runtime efficiency either via caching scene reference poses in a hierarchical tree structure to fasten the searching process or leveraging any available prior information such as GPS coordinates to quickly find a subset of poses.
Indoor performance bottleneck.  NeRF predicted depth maps are used to compute pseudo ground-truth for matching supervision. Incorrect depth predictions can lead to misaligned feature correspondences. In contrast, image matching, SCR, and APR methods use more accurate labels like Colmap camera poses or 3D maps. For small-scale indoor scenes, precise supervision is essential to achieve centimeter-level errors. Our method based on feature matching, however, scales better than regression-based approaches in larger outdoor scenes. Introducing uncertainty measures to ignore inaccurate matches, as in  [15], and improved NeRF reconstructions with accurate depth maps will benefit our method.